ihave a problem.i needd help!i just cannot fiqure out this word problem about age.well hope you can help. here itis
Sam is 10 years older than Jack.Twice Sam's age subtracted from 6 times jack's age is 7 more than Jack's age. find both sam and jack's age.
i cant even say it i get a headache i dont know where to start help me plz

Answer by stanbon(75887)

Sam is 10 years older than Jack.Twice Sam's age subtracted from 6 times jack's age is 7 more than Jack's age. find both sam and jack's age.
---------------
Let Jack's age be "x"; Sam's age is "x+10"
EQUATION:
6x-2(x+10)= x+7
6x-2x-20 = x + 7
3x = 27
x = 9 (Jack's age
x+10 = 19 (Sam's age)
=============
Cheers,
Stan H.